OpenStudy (anonymous):

solve for the indicated letter a = 5b the solution is b =

OpenStudy (anonymous):

b=a/5

OpenStudy (anonymous):

\[\frac{a}{5}\]

OpenStudy (katrinakaif):

Divide by 5 on both sides to get b alone so a/5 = b